OpenStudy (anonymous):

Find the slope of the line passing through (11, -2) and (2, -2). A. 4/9 B. 9 C. 0 D. Undefined

jimthompson5910 (jim_thompson5910):

Two generic points that define a straight line are (x1, y1) and (x2, y2) In this case, (x1, y1) = (11, -2), which means x1=11 and y1= -2 and (x2, y2) = (2, -2), which means x2=2 and y2= -2 So x1=11, y1= -2, x2=2, and y2= -2. ------------------- Now use the slope formula m = (y2-y1)/(x2-x1) m = ( -2-( -2))/(2-11) m = ( -2+2)/(2-11) m = 0/(-9) m = 0 So the slope of the line that passes through the two points is 0